# ROSE: A Recognition-Oriented Speech Enhancement Framework in Air Traffic Control Using Multi-Objective Learning

### v1

@@ Introduction
Sequence Path. This path is to guide the model to focus on selective frames by employing higher attention weights. Specifically, the trainable attention module focuses on the marginal frames to remove echo features to support the SE task and also highlights the central frames of each phoneme to advance the ASR task.
Fusion path. This path is to fuse features learned from channel and sequence paths and feeds them into subsequent modules.
The global average pooling (GAP) operation is applied to generate the sample-related initial weights based on the input feature map. As shown in Figure 2 (c), let the input feature map be $X\in\mathbb{R}^{C\times L}$, where $C$ is the number of channels, $L$ is the length of the sequence dimension. The GAP operation is applied to generate the sample-related initial weights based on the input feature map along the channel dimension. The squeeze-and-excitation attention is to obtain the channel attention weights $W_{C}$, as shown below:
$BX=\delta\left(\psi_{\frac{C}{r}}^{[1\times 1]}(GAP(X))\right),BX\in\mathbb{R}^{\frac{C}{r}\times 1}$ (3)
$W_{C}=\sigma\left(\psi_{C}^{[1\times 1]}(BX)\right),W_{C}\in\mathbb{R}^{C\times 1}$ (4)
where Equation 3 and 4 denote the channel squeeze and excitation operations, respectively. $r$ is a hyper-parameter to indicate the squeezed degree. $\delta$ denotes the ReLU activation. Finally, the output feature of the channel attention $X_{C}$ is obtained by the following rules:
@@
Similarly, the sequence attention weight $W_{L}$ is obtained by using Conv1d of 1 filter to convolve the input feature $X$ with the activation function. The sequence-attentive feature $X_{L}$ is obtained by Equation 7:
$W_{L}=\sigma\left(\psi_{1}^{[1\times 1]}(X)\right),W_{L}\in\mathbb{R}^{1\times L}$ (6)
@@
Finally, The fused features of the CSAtt block are achieved as below:
@@
Multi-objective Loss Function
As illustrated before, the proposed model is optimized by multi-objective learning to enhance speech quality and ASR-required features. The MAE loss in the time domain and log-scale STFT-magnitude loss in the frequency domain serve as the optimization targets for the SE task.
@@
$L_{MAG}=\left\|\log\left|STFT\left(s;\theta\right)\right|-\log\left|STFT\left(\tilde{s};\theta\right)\right|\right\|_{F}$ (10)
@@
where *$\left\|*\right\|_{1}$ and *$\left\|*\right\|_{F}$ denote the $L_{1}$ and Frobenius norm, respectively. $\left|STFT(*)\right|$ denotes the magnitude of signal processed through STFT. $\theta$ denotes the STFT parameters (FFT bins, hop sizes, window lengths).
For the ASR task, the spectral convergence (SC) loss is regarded as the optimization target to achieve ASR-required features. Considering that the common ASR features are based on the STFT operation, such as spectrogram and MFCC. According to (Yamamoto, Song, and Kim 2020), the SC loss can be defined as follows:
@@
where $D$ denotes the handcrafted features of the speech signal. In this work, spectrogram and MFCC features are selected to implement ASR task $L_{ASR}$, as shown below:
$L_{Spec}=f\left(Spectrogram;\theta\right)\hskip 5.69046ptL_{MFCC}=f\left(MFCC;\theta\right)$ (13)
$L_{ASR}=L_{Spec}+L_{MFCC}$ (14)
@@
Experiments
Dataset
A real-world dataset is collected from an industrial ATC system. Detailed descriptions of dataset are shown in Appendix. As mentioned above, the ATCO speech with echo is defined as noisy and corresponding echo-free speech is defined as clean. All samples are multilingual with various ATC instructions and are sampled at 8 kHz. The text labels are annotated for the test set to evaluate the ASR performance.
Experimental Configurations
In the ROSE, the depth of both the encoder and decoder are 5 and the hidden dimension of each layer is $H=48$. For the CNN block of encoder/decoder, Conv1d or ConvTr1d is configured with kernel size $K=8$ and stride $S=4$. For the CSAtt block, $r$ is set to 2 for squeezing the feature channels. Two bidirectional LSTM layers with 768 neurons are designed to capture the sequential dependencies. For the STFT parameters, FFT bins are 512, hop sizes are 100, and window lengths are 400. The dimension of MFCC is set to 13. The $\lambda_{i}$ for the loss function in Equation 15 is set to 1. The Adam optimizer is applied to optimize the proposed framework with an initial learning rate of 3e-4 and 0.999 decay. All samples are upsampled from 8 kHz to 16 kHz and clipped to 4 seconds long. The batch size is set to 64 during the model training.
To confirm the effectiveness and efficiency of the proposed model on ASR performance, the test set is evaluated using a Deepspeech 2 (DS2) (Amodei et al. 2016) based model trained on the ATCSpeech corpus (Yang et al. 2020).
@@ 4 Can understand ATC instructions clearly and repeat the spoken words
Table 1: Details of the ATC-related subjective measure.
Baseline Methods and Evaluation Metrics
In this work, several open-source SOTA baselines with the different domains are selected to validate the proposed model, which are classified as follows: 1) time domain: Wiener (Lim and Oppenheim 1978), SEGAN (Pascual, Bonafonte, and Serra 2017), DEMUCS (Defossez, Synnaeve, and Adi 2020) and SE-Conformer (Kim and Seo 2021), 2) T-F domain: CRN (Tan and Wang 2019), DCCRN-E (Hu et al. 2020), MetricGAN+ (Fu et al. 2021) and FullSubNet+ (Chen et al. 2022). All of those methods are conducted on the official configurations.
For the metrics of the SE task, we compute the following five typical objective measures on the test set, i.e., mean opinion score (MOS) prediction of the $i)$ signal distortion attending only to the speech signal (CSIG), $ii)$ intrusiveness of background noise (CBAK), $iii)$ the overall effect (COVL) (from 1 to 5) (Hu and Loizou 2008), perceptual evaluation of speech quality (PESQ) (from -0.5 to 4.5) (Recommendation 2001), short-time objective intelligibility (STOI) (from 0 to 100) (Taal et al. 2011).
For the subjective evaluation, we randomly sample 10 utterances with complete ATC instructions and each one was scored by 45 ATC experts along two facets: 1) a relevant subjective MOS study is recommended in ITU-T P.835 (recommendation 2003), including level of signal distortion (SIG), intrusiveness of background noise (BAK), and overall quality (OVL). 2) an ATC-related subjective measure is proposed to rate the ATC instructions recognition (AIR) and detailed descriptions are shown in Table 1. For all the above metrics, a larger score indicates a higher speech quality.
For the metric of the ASR task, the label error rate (LER) based on the minimum basic unit of languages is applied to evaluate the ASR performance.
@@
A1 Noisy 3.28 1.81 2.48 1.80 56.39 3.42 2.84 2.91 3.4 4.66
A2 Wiener (T) 2.17 1.41 1.66 1.49 54.20 2.78 2.13 2.26 3.1 10.62
A3 SEGAN (T) 3.56 2.55 2.79 2.12 69.90 3.47 2.85 2.93 3.4 9.74
A4 CRN (T-F) 3.57 2.51 2.89 2.33 70.91 3.54 2.78 2.94 3.5 7.97
A5 DEMUCS (T) 4.25 2.95 3.45 2.66 80.34 3.89 3.21 3.15 3.6 4.72
A6 DCCRN-E (T-F) 3.76 2.69 3.07 2.45 70.64 3.62 3.03 2.95 3.5 8.46
A7 MetricGAN+ (T-F) 4.13 3.12 3.41 2.97 72.19 3.78 3.19 3.02 3.5 7.98
A8 SE-Conformer (T) 4.40 3.15 3.69 3.02 81.23 3.83 3.21 3.13 3.7 4.28
A9 FullSubNet+ (T-F) 4.12 3.00 3.45 2.82 74.85 3.82 3.18 3.10 3.6 5.35
A10 ROSE (Ours) (T) 4.58 3.22 3.85 3.09 83.90 3.98 3.24 3.15 3.8 3.50
B1 A10 $w/o$ $L_{Spec}$ 4.42 3.08 3.64 2.85 83.04 - - - - 4.00
B2 A10 $w/o$ $L_{MFCC}$ 4.40 3.04 3.64 2.88 82.65 - - - - 4.08
B3 A10 $w/o$ $L_{ASR}$ 4.33 3.01 3.53 2.71 82.15 - - - - 4.29
C1 B3 $w/o$ ABSF 4.27 2.94 3.46 2.66 81.60 - - - - 4.31
C2 B3 $w/o$ CSAtt 4.15 2.84 3.33 2.51 79.79 - - - - 4.36
C3 B3 $w/o$ $attn.$ 4.04 2.76 3.21 2.39 79.37 - - - - 4.40
Table 2: Experiments for different methods and ablation studies on ATC corpus. The LER of clean signal is 2.25.
@@
Noisy - 3.35 2.44 2.63 1.97 92
Wiener (Lim and Oppenheim 1978) Waveform 3.23 2.68 2.67 2.22 93
SEGAN (Pascual, Bonafonte, and Serra 2017) Waveform 3.48 2.94 2.80 2.16 -
CRN (Tan and Wang 2019) Magnitude 3.59 3.11 2.71 2.48 93
MetricGAN (Fu et al. 2019) Magnitude 3.99 3.18 3.42 2.86 -
CRGAN (Zhang et al. 2020) Magnitude 4.16 3.24 3.54 2.92 94
DEMUCS (Defossez, Synnaeve, and Adi 2020) Waveform 4.31 3.40 3.63 3.07 95
DCCRN-E (Hu et al. 2020) RI 3.74 3.13 2.75 2.54 94
MetricGAN+ (Fu et al. 2021) Magnitude 4.14 3.16 3.64 3.15 -
TSTNN (Wang, He, and Zhu 2021) Waveform 4.17 3.53 3.49 2.96 95
SE-Conformer (Kim and Seo 2021) Waveform 4.45 3.55 3.82 3.13 95
CTS-Net (Li et al. 2021) RI+Magnitude 4.25 3.46 3.59 2.92 -
FullSubNet+ (Chen et al. 2022) RI+Magnitude 3.86 3.42 3.57 2.88 94
GaGNet (Li et al. 2022) RI+Magnitude 4.26 3.45 3.59 2.94 95
DeepFilterNet2 (Schröter et al. 2022) RI+Magnitude 4.30 3.40 3.70 3.08 94
ROSE (Ours) Waveform 4.47 3.56 3.72 3.01 95
Table 3: SE results on Voice Bank + DEMAND dataset, “-” denotes the result is not provided in the associated papers.
Results and Discussions
Performance Evaluation
The experimental results for the performance comparison are shown in Table 2 (group A). The following conclusions can be obtained from the experimental results:
(1) From experiments A1 and A2, the traditional SE method (Wiener) fails to suppress the speech echo, and even causes extra noise to deteriorate the quality of the original speeches. The results also indicate that the speech echo is a dedicated phenomenon in the ATC domain, which cannot be addressed by the traditional denoising method.
(2) According to experiments A3 and A7, although the GAN-based method improves the objective and subjective metrics of the noisy speech, it deteriorates the ASR accuracy due to the generation mechanism for pseudo-speech by learning the data distribution of clean signals. The main purpose is to satisfy the auditory requirements of the speech quality but fails to enclose ASR-required features to achieve desired recognition accuracy.
(3) According to the supervised-based DNN models, the time-domain methods (A5, A8 and A10) are generally superior to the T-F domain methods (A4, A6, A7 and A9) in terms of objective metrics, and have a significant advantage in ASR performance. The results show that the T-F methods use T-F representatives operated by STFT as input, which inevitably discards potential acoustic and linguistic features, leading to undesirable results.
(4) By focusing on the ASR performance, all the comparative baselines suffer inferior LER compared to the noisy speeches (only marginal improvements obtained by the A8) since they are SE-oriented approaches to improve the auditory quality. The metrics of AIR indicate that improving the clarity and intelligibility of speech can indeed improve the rater’s ability to understand ATC instructions.
Among all comparative methods, the proposed model ROSE achieves the best performance for both the SE and ASR tasks.
@@
Mixed - 33.36 42.72
CRN Magnitude 35.68 38.23
SE-Conformer Waveform 21.43 31.10
DCCRN-E RI 26.69 33.67
ROSE (Ours) Waveform 12.75 22.31
$w/o$ $L_{ASR}$ Waveform 19.12 29.44
$w/o$ $attn.$ Waveform 17.86 26.56
test-clean - 9.97
Table 4: ASR results on two artificial datasets. ”$a$” denotes additive noise, ”$e$” denotes simulative echo.
Ablation Study

### v2

@@ III-D The Channel and Sequence Attention
Sequence Path. This path is to guide the model to focus on selective frames by employing higher attention weights. Specifically, the trainable attention maps focus on the marginal frames to remove echo features to support the SE task and also highlight the central frames of each phoneme to advance the ASR task.
Fusion path. This path is to fuse features learned from channel and sequence paths and feed them into subsequent modules.
The global average pooling (GAP) operation is applied to generate the sample-related initial weights based on the input feature map. As shown in Figure 2 (c), let the input feature map be $X\in\mathbb{R}^{C\times L}$, where $C$ is the number of channels, $L$ is the length of the sequence dimension. The GAP operation is applied to generate the sample-related initial weights based on the input feature map along the channel dimension. The squeeze-and-excitation attention [41] is to obtain the channel attention weights $W_{C}$, as shown below:
$X_{B}=\delta\left(\psi_{\frac{C}{r}}^{[1,1]}(GAP(X))\right),X_{B}\in\mathbb{R}^{\frac{C}{r}\times 1}$ (5)
$W_{C}=\sigma\left(\psi_{C}^{[1,1]}(X_{B})\right),W_{C}\in\mathbb{R}^{C\times 1}$ (6)
where Equations 5 and 6 denote the channel squeeze and excitation operations, in which the 1-d convolution layer followed by the ReLU activation is used to squeeze the number of channels and the 1-d convolution layer followed by the sigmoid activation is used to excite the number of channels, respectively. $r$ is a hyper-parameter to indicate the squeezed degree. $\delta$ denotes the ReLU activation.
Finally, the output feature of the channel attention $X_{C}$ is obtained by the following rules:
@@
Similarly, a 1-d convolution layer followed by a sigmoid activation is applied to reduce the number of channels to 1, obtaining the sequence attention weights $W_{L}$. The sequence-attentive feature $X_{L}$ is obtained by Equation 9:
$W_{L}=\sigma\left(\psi_{1}^{[1,1]}(X)\right),W_{L}\in\mathbb{R}^{1\times L}$ (8)
@@
Finally, The fused features of the CSAtt module are achieved as below:
@@
III-E Multi-objective Loss Function
As illustrated before, the proposed model is optimized by multi-objective learning to enhance speech quality and ASR-required features. Based on previous works [12, 13], the MAE loss in the time domain and log-scale STFT-magnitude loss in the frequency domain serve as the optimization targets for the SE task.
@@ III-E Multi-objective Loss Function
$L_{MAG}=\left\|\log\left|STFT\left(s;\theta\right)\right|-\log\left|STFT\left(\tilde{s};\theta\right)\right|\right\|_{1}$ (12)
@@
where $\left\|*\right\|_{1}$ denotes the $L_{1}$ norm. $\left|STFT(*)\right|$ denotes the magnitude of signal processed through STFT. $\theta$ denotes the STFT parameters (FFT bins, hop sizes, window lengths).
For the ASR task, the spectral convergence (SC) loss is regarded as the optimization target to achieve ASR-required features. According to [42], the core idea is to optimize the spectral distance between the enhanced speech and the reference speech, aiming at guiding the model to capture the informative ASR-related information embedded in the signal spectrums. The SC loss can be defined as follows:
@@
where $\left\|*\right\|_{F}$ denotes the Frobenius norm and $D$ denotes the handcrafted features of the speech signal. In this work, as the input features of the existing ASR models in the ATC domain [2, 15, 4], spectrogram and MFCC features are selected to achieve ASR-oriented loss $L_{ASR}$, as shown below:
$L_{Spec.}=f\left(\text{Spectrogram};\theta\right),\hskip 5.69046ptL_{MFCC}=f\left(\text{MFCC};\theta\right)$ (15)
$L_{ASR}=L_{Spec.}+L_{MFCC}$ (16)
@@
IV Experiments
IV-A Dataset
In this work, three datasets collected from real-world industrial ATC systems are used to validate the proposed approach, including clean-noisy pair dataset [6], ATCSpeech-large dataset [17] and SSL dataset [43]. Specifically, The clean-noisy pair dataset contains only ATCO’s speech data and is used to train and test proposed ROSE and other baseline methods. The detailed descriptions of dataset division are shown in Table I and more relevant details can be found in [6]. In addition, the text labels are annotated for the dataset to evaluate the ASR performance, including a total of 683 Chinese characters and 437 English words (a total of 712 tokens in the vocabulary, including 683 Chinese characters, 26 English letters and 3 special tokens, i.e., SPACE, UNK and ’). The ATCspeech-large dataset contains a total of 1275 hours of annotated speech data, including 701 Chinese characters and 1221 English words (a total of 730 tokens in the vocabulary). The SSL dataset contains two sub-datasets for pre-training (SSL) and fine-tuning (SL) processes, respectively, in which the SSL set contains a total of 1032 hours of unlabeled speech data and the SL set contains a total of 15 hours of annotated speech data, including a total of 639 Chinese characters and 584 English words (a total of 668 tokens in the vocabulary). All datasets are independent and the latter two datasets are used separately to train the two selected ASR models.
TABLE I: Data size of the dataset, #U denotes the speech utterances and #H denotes the speech durations (Hours).
Language Train Validation Test
#U #H #U #H #U #H
Chinese 42189 37.28 4188 3.69 6012 5.08
English 5064 5.55 558 0.62 502 0.54
Total 47253 42.83 4746 4.31 6514 5.62
IV-B Experimental Configurations
In the ROSE, the depth of both the encoder and decoder are 5 and the hidden dimension of each layer is $H=48$. For the basic CNN block of encoder/decoder, Conv1d or ConvTr1d is configured with kernel size $K=8$ and stride $S=4$. For the CSAtt module, $r$ is set to 2 for squeezing the feature channels. Two bidirectional LSTM layers with 768 neurons are designed to capture the sequential dependencies. For the STFT parameters, FFT bins are 512, hop sizes are 100, and window lengths are 400. The dimension of MFCC is set to 13. The $\lambda_{i}$ for the loss function in Equation 17 is set to 1. The Adam optimizer is applied to optimize the proposed framework with an initial learning rate of 3e-4 and 0.999 decay. All samples in the clean-noisy pair dataset are upsampled from 8 kHz to 16 kHz and clipped to 4 seconds long for fair comparisons with baseline models. The batch size is set to 64 during the model training.
To confirm the effectiveness and efficiency of the proposed model on ASR performance, two optimized ASR models are applied to evaluate the clean-noisy pair test set, including a supervised Deepspeech 2 (DS2) model [44] and a self-supervised learning (SSL)-based Wav2Vec 2.0 (W2V2) model [43]. Specifically, 1:) the ATCSpeech-large dataset is applied to train the DS2 model, in which the input feature is the 81-dimension spectrogram. The Adam optimizer is applied to optimize the model parameters with an initial learning rate of 1e-3 and 0.99 decay. 2:) The SSL set is applied to pre-train the W2V2 encoder to obtain a pre-trained model, and the SL set is applied to supervisedly fine-tune the pre-trained W2V2 model to obtain an ASR-focused fine-tuned model. The W2V2 encoder is optimized with Adam and trained for 450k updates, warming up the learning rate for the first 8% of updates to a peak of 5 ×$\times 10^{-4}$, and then linearly decaying it. The finetuning process is implemented on fairseq [^1] https://github.com/pytorch/fairseq/tree/master/examples/wav2vec with the Connectionist Temporal Classification (CTC) loss function. All the above ASR models are end-to-end multilingual models without language models. Note that the ASR experiments designed in this work focus on validating the proposed ROSE on both SE and ASR tasks in the ATC domain.
TABLE II: Details of the AIR subjective rating.
@@ 4 Can understand ATC instructions clearly and repeat the spoken words
IV-C Baseline Methods
In this work, several open-source SOTA baselines with different types are selected to validate the proposed model, which are classified as follows: 1) traditional methods: Wiener [28] and spectral subtraction (Spec-Sub)[29], 2) GAN-based generative methods: SEGAN [39], MetricGAN [30] and MetricGAN+ [31], 3) Diffusion-based generative methods: DiffuSE [45], CDiffuSE [46] and SGMSE [47] and 4) discriminative methods: CRN [48], DCCRN-E [35], DEMUCS [12], BSRNN [32], SE-Conformer [13], FullSubNet+ [49], DB-AIAT [37] and TF-GridNet [38]. These methods cover time (T) and T-F domain models, as well as causal and non-causal paradigms. All of those methods are conducted on the official configurations and adapted to the ATC corpus following the relevant works.
IV-D Evaluation Metrics
For the metrics of the SE task, we compute the following five typical objective measures on the test set, i.e., mean opinion score (MOS) prediction of the $i)$ signal distortion attending only to the speech signal (CSIG), $ii)$ intrusiveness of background noise (CBAK), $iii)$ the overall effect (COVL) (from 1 to 5) [50], perceptual evaluation of speech quality (PESQ) (from -0.5 to 4.5) [51], short-time objective intelligibility (STOI) (from 0 to 1) [52].
For the subjective evaluation, we randomly sample 10 utterances with complete ATC instructions and each one was scored by 45 ATC experts along two facets: 1) a relevant subjective MOS study is recommended in ITU-T P.835 [53], including level of signal distortion (SIG), intrusiveness of background noise (BAK), and overall quality (OVL). 2) an ATC-related subjective measure is proposed to rate the ATC instructions recognition (AIR) and detailed descriptions are shown in Table II. For all the above metrics, a larger score indicates a higher speech quality.
For the metric of the ASR task, the CER based on the Chinese character and English letter is applied to evaluate the ASR performance. We report the mean results and their 95% confidence interval. For each model, the number of model parameters is reported in millions (M). We use the calflops [^2] https://github.com/MrYxJ/calculate-flops.pytorch toolkit to count the number of multiply-accumulate operations (MAC) for processing a 4-second signal, and report it in giga-operations per second (GMAC/s).
TABLE III: Objectvie experimental results for different methods on ATC corpus. The CER of the DS2 and W2V2 of the clean signal are (2.25$\pm$.16)and (1.37$\pm$.15)%, respectively.
[…]
(M) DS2 W2V2
Noisy - - - - 3.28$\pm$.01 1.81$\pm$.01 2.48$\pm$.01 1.80$\pm$.01 0.56$\pm$.00 4.66$\pm$.24 2.81$\pm$.17
Traditional Methods
@@ IV-D Evaluation Metrics
GAN-based Generative Methods
@@
[…]
Diffusion-based Generative Methods
[…]
Discriminative Methods
CRN [48] T-F 1.70 8.80 3.57$\pm$.04 2.51$\pm$.04 2.89$\pm$.04 2.33$\pm$.04 0.71$\pm$.00 7.97$\pm$.31 2.91$\pm$.25
DCCRN-E[35] T-F 3.67 33.48 3.76$\pm$.04 2.69$\pm$.03 3.07$\pm$.04 2.45$\pm$.03 0.71$\pm$.00 8.46$\pm$.34 3.97$\pm$.29
DEMUCS [12] T 53.52 37.83 4.22$\pm$.04 2.91$\pm$.04 3.41$\pm$.04 2.62$\pm$.04 0.79$\pm$.01 4.89$\pm$.21 2.75$\pm$.21
BSRNN [32] T-F 2.96 2.57 4.55$\pm$.04 3.26$\pm$.04 3.85$\pm$.04 3.04$\pm$.04 0.83$\pm$.00 5.66$\pm$.24 2.86$\pm$.29
[…]
TABLE IV: Subjective results for different methods on selected ATC samples. The selected DEMUCS is non-causal model.
[…]
Noisy 3.42$\pm$.10 2.84$\pm$.09 2.91$\pm$.08 3.41$\pm$.03
Wiener[28] 2.78$\pm$.12 2.13$\pm$.11 2.26$\pm$.10 3.13$\pm$.04
Spec-Sub[29] 2.98$\pm$.09 2.27$\pm$.10 2.39$\pm$.11 3.23$\pm$.03
SEGAN[39] 3.47$\pm$.10 2.85$\pm$.10 2.93$\pm$.09 3.43$\pm$.04
MetricGAN[30] 3.76$\pm$.11 3.21$\pm$.09 3.03$\pm$.10 3.43$\pm$.03
MetricGAN+ [31] 3.78$\pm$.10 3.19$\pm$.09 3.02$\pm$.09 3.51$\pm$.04
DiffuSE [45] 3.57$\pm$.10 2.84$\pm$.10 2.97$\pm$.10 3.41$\pm$.03
CDiffuSE [46] 3.55$\pm$.09 2.88$\pm$.10 2.96$\pm$.10 3.51$\pm$.03
SGMSE [47] 3.62$\pm$.10 2.97$\pm$.10 3.02$\pm$.08 3.62$\pm$.03
CRN [48] 3.54$\pm$.08 2.78$\pm$.09 2.94$\pm$.10 3.47$\pm$.03
DCCRN-E [35] 3.62$\pm$.10 3.03$\pm$.10 2.95$\pm$.10 3.52$\pm$.03
DEMUCS [12] 3.89$\pm$.10 3.21$\pm$.10 3.15$\pm$.10 3.56$\pm$.04
BSRNN [32] 3.85$\pm$.09 3.23$\pm$.10 3.13$\pm$.10 3.69$\pm$.03
SE-Conformer [13] 3.83$\pm$.10 3.21$\pm$.10 3.13$\pm$.10 3.68$\pm$.03
FullSubNet+ [49] 3.82$\pm$.09 3.18$\pm$.08 3.10$\pm$.09 3.60$\pm$.03
DB-AIAT [37] 3.91$\pm$.10 3.20$\pm$.10 3.16$\pm$.09 3.73$\pm$.03
TF-GridNet [38] 3.94$\pm$.10 3.25$\pm$.10 3.10$\pm$.10 3.74$\pm$.04
ROSE (Ours) 3.98$\pm$.09 3.24$\pm$.10 3.15$\pm$.10 3.78$\pm$.03
TABLE V: The experimental results of the ablation study.
[…]
(M) DS2 W2V2
A1 ROSE 36.98 4.73 4.58$\pm$.04 3.22$\pm$.04 3.85$\pm$.03 3.09$\pm$.03 0.84$\pm$.00 3.50$\pm$.18 2.42$\pm$.19
B1 A1 $w/o$ $L_{MAE}$ 36.98 4.73 4.32$\pm$.03 2.99$\pm$.04 3.51$\pm$.03 2.73$\pm$.04 0.82$\pm$.00 3.88$\pm$.20 2.45$\pm$.19
B2 A1 $w/o$ $L_{MAG}$ 36.98 4.73 4.36$\pm$.03 3.03$\pm$.04 3.54$\pm$.04 2.75$\pm$.04 0.82$\pm$.00 3.86$\pm$.20 2.47$\pm$.20
B3 A1 $w/o$ $L_{SE}$ 36.98 4.73 4.28$\pm$.04 2.96$\pm$.04 3.48$\pm$.04 2.69$\pm$.04 0.81$\pm$.00 3.91$\pm$.22 2.50$\pm$.19
B4 A1 $w/o$ $L_{Spec.}$ 36.98 4.73 4.42$\pm$.03 3.08$\pm$.04 3.64$\pm$.04 2.85$\pm$.04 0.83$\pm$.00 4.00$\pm$.19 2.53$\pm$.20
B5 A1 $w/o$ $L_{MFCC}$ 36.98 4.73 4.40$\pm$.04 3.04$\pm$.04 3.64$\pm$.04 2.88$\pm$.04 0.83$\pm$.00 4.08$\pm$.21 2.56$\pm$.23
B6 A1 $w/o$ $L_{ASR}$ 36.98 4.73 4.33$\pm$.04 3.01$\pm$.03 3.53$\pm$.04 2.71$\pm$.04 0.82$\pm$.00 4.29$\pm$.24 2.60$\pm$.24
C1 B6 $w/o$ ABSF 35.80 4.18 4.27$\pm$.04 2.94$\pm$.04 3.46$\pm$.04 2.66$\pm$.04 0.82$\pm$.00 4.31$\pm$.24 2.63$\pm$.24
C2 B6 $w/o$ CSAtt 35.40 4.72 4.15$\pm$.03 2.84$\pm$.04 3.33$\pm$.04 2.51$\pm$.04 0.80$\pm$.00 4.36$\pm$.24 2.68$\pm$.28
C3 B6 $w/o$ $attn.$ 34.22 4.17 4.04$\pm$.04 2.76$\pm$.04 3.21$\pm$.04 2.39$\pm$.03 0.79$\pm$.00 4.40$\pm$.26 2.71$\pm$.27
TABLE VI: SE results on Voice Bank + DEMAND dataset. “—” denotes the results are not provided in the associated papers. “*” denotes that the results are reproduced by us. Optimal values are bold and sub-optimal values are marked with underlined wavy line in each column.
[…]
Noisy - - - 3.35 2.44 2.63 1.97 0.92
Traditional Methods
Wiener [28] Waveform - - 3.23 2.68 2.67 2.22 0.93
Spec-Sub [29] * Magnitude - - 3.40 2.71 2.74 2.26 0.93
GAN-based Generative Methods
SEGAN [39] Waveform 65.82 13.39 3.48 2.94 2.80 2.16
MetricGAN [30] Magnitude 1.86 28.51 3.99 3.18 3.42 2.86
CRGAN [54] Magnitude 4.16 3.24 3.54 2.92 0.94
MetricGAN+ [31] Magnitude 1.86 28.51 4.14 3.16 3.64 3.15
MetricGAN-OKD (PE+CS) [55] Magnitude 0.82 4.23 3.07 3.73 3.24
Diffusion-based Generative Methods
DiffuSE [45] * Magnitude 2.23 210.31 3.64 2.85 3.02 2.43 0.90
CDiffuSE [46] * Magnitude 4.28 190.23 3.85 3.03 3.20 2.45 0.91
SGMSE [47] * Magnitude 3.89 152.77 4.30 3.52 3.67 2.93 0.94
Discriminative Methods
CRN [48] Magnitude 1.70 8.80 3.59 3.11 2.71 2.48 0.93
DEMUCS [12] Waveform 68.87 38.12 4.31 3.40 3.63 3.07 0.95
DCCRN-E [35] RI 3.67 33.48 3.74 3.13 2.75 2.54 0.94
PHASEN [36] Phase+Magnitude 4.21 3.55 3.62 2.99
TSTNN [33] Waveform 0.92 4.17 3.53 3.49 2.96 0.95
SE-Conformer [13] Waveform 75.28 44.96 4.45 3.55 3.82 3.13 0.95
DB-AIAT [37] * RI+Magnitude 2.81 41.8 4.41 3.53 3.76 3.21 0.95
FullSubNet+ [49] RI+Magnitude 8.67 48.57 3.86 3.42 3.57 2.88 0.94
GaGNet [56] RI+Magnitude 4.26 3.45 3.59 2.94 0.95
DeepFilterNet2 [57] RI+Magnitude 2.31 4.30 3.40 3.70 3.08 0.94
MANNER-S-8.1GF [40] Waveform 1.38 4.09 4.40 3.55 3.74 3.01 0.95
BSRNN [32] * Magnitude 2.96 2.57 4.41 3.59 3.80 3.05 0.95
TF-GridNet [38] * RI 8.2 19.2 4.37 3.50 3.74 2.98 0.94
ROSE (Ours) Waveform 36.98 4.73 4.47 3.56 3.72 3.01 0.95
V Results and Discussions
V-A Performance Evaluation
The objective and subjective experimental results for the performance comparison are shown in Table III and Table IV, respectively. The following conclusions can be obtained from the experimental results:
1) The traditional SE methods (Wiener and Spec-Sub) fail to suppress the speech echo, and even cause extra noise to deteriorate the quality of the original speeches. The results also indicate that the speech echo is a dedicated phenomenon in the ATC domain, which cannot be addressed by the traditional denoising method.
2) Although the GAN-based and Diffusion-based generative methods improve the objective and subjective SE-related metrics of the noisy speech, they deteriorate the ASR accuracy due to the generation mechanism for pseudo-speech by learning the data distribution of clean signals. The main purpose is to satisfy the auditory requirements of the speech quality but fails to enclose ASR-required features to achieve desired recognition accuracy.
3) According to the discriminative models, although the time-domain methods are slightly inferior to the T-F methods in terms of SE-related metrics, they have a significant advantage in ASR performance. The results show that the original waveform contains the complete information of the speech, while the input features (e.g., magnitude, phase, complex and their combinations) of the T-F methods may fail to present the same complete features as the waveform, further leading to the lack of critical ASR-related linguistic information, especially in ATC scenarios.
4) Compared to the causal models, the non-causal models obtain superior performance on both SE and ASR tasks. In general, the non-causal paradigm can leverage the sequential information in the past and future dimensions to capture the long-distance temporal dependencies among the signal frames.
5) By focusing on the ASR performance, almost all the comparative baselines suffer inferior CER compared to the noisy speeches (only marginal improvements in certain models) since they are SE-oriented approaches to improve the auditory quality. The metrics of AIR indicate that improving the clarity and intelligibility of speech can indeed improve the rater’s ability to understand ATC instructions.
In summary, the proposed approach aims to address the ATC speech echo and achieves the desired performance improvement on the real-world dataset. Although the proposed ROSE only achieves sub-optimal computational costs, fortunately, it generally achieves the best performance in ATC scenarios, which is the preferred goal of this work. Furthermore, the proposed model achieves improved recognition performance both for the feature-adapted DS2 model and the SSL-based W2V2 model, which further confirms the generalization of the proposed technical improvements.
V-B Ablation Study
[…]